Solve the equation by completing the square
x^2 - 14x + 24 = 0

Answer by scott8148(6628)

x^2-14x=-24 ... half of -14 is -7, (-7)^2 is 49

x^2-14x+49=-24+49 ... (x-7)^2=25 ... x-7=+-5

x=-2, x=-12